OpenStudy (anonymous):

((sinx)^4/(tanx)^2 ) Simplify.

OpenStudy (zzr0ck3r):

=\[\frac{sinx^{4}}{(\frac{sinx^2}{cosx^2})}=\frac{sinx^{4}cosx^{2}}{sinx^{2}}\]

OpenStudy (anonymous):

= sin^2cos^2

OpenStudy (mayankdevnani):

when we simplify the @zzr0ck3r answer: we get \[\huge \frac{sinx^4cosx^2}{sinx^2}\] \[\huge = sinx^2cosx^2\]

OpenStudy (mayankdevnani):

ok... @TabletofDaFuta

OpenStudy (anonymous):

o thats the most simplified...? i got it right then lololol but thx

OpenStudy (mayankdevnani):

welcome:)